OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, 7) and perpendicular to y =-1/4 -x - 5. y = -4x - 11 y = -4x - 3 y = 4x + 11 y = 4x + 3

OpenStudy (anonymous):

a perpendicular line has the "negative inverse" slope of the original line. So for example, a line with slope 2 has a perp line with slope -(1/2) Find the slope of the original line, then take the neg inverse to get the perp line slope. That's the first part of the solution. Then you need to use the point (-1,7) and write the equation of the perp line through that point using the slope and that point... it's called point-slope form, and it looks like y - y1 = m(x - x1) where m is the slope and your point (-1,7) is (x1,y1). Then simplify and move the terms around to end up with the y = mx + b form like the answer options show. Sounds like a lot, but it's just a couple of steps....